OpenStudy (anonymous):

help!...how :given that 4t=18-2t,find the value of t

sam (.sam.):

4t=18-2t Move all terms not containing t to the right-hand side of the equation. 6t=18 Divide each term in the equation by 6. t=3

OpenStudy (anonymous):

4T=18-2t 6t=18 t=3

OpenStudy (amistre64):